I need help with this problem....
Problem :
100(1.04)^2t=5000
I dont understand how to solve for t in terms of logarithms ... could someone please help.. Thanks

Answer by RAY100(1637)

100(1.04)^2t=5000
(1.04)^2t=50
log both sides
log((1.04)^2t)=log50
but logx^y=ylogx
2tlog(1.04)=log50
2t=log50/log(1.04)=99.743
divide by 2
t=49.87
check
100(1.04)^2(49.87)=5000
100(1.04)^99.743=5000
5000=5000,,,ok